OpenStudy (anonymous):

integrate (3x+2)/(x+1)

OpenStudy (anonymous):

reform it to: \[\frac{ 3x + 2 }{ x+1 } = \frac{ 3(x+1) - 1 }{ x+1 } = 3\frac{ x+1 }{ x+1 } - \frac{ 1 }{ x+1 } = 3 - \frac{ 1 }{ x+1 }\] i'm sure you can integrate this ^_^

OpenStudy (anonymous):

Thank you very much ^_^